A recent study at a local college claimed that the proportion, p, of students who commute more than fifteen miles to school ianswer neatly and correctly please!
Answer #1

Solution :

Given that,

p_{0} = 0.20

1 - p_{0} = 0.80

n = 270

x = 58

Level of significance = \alpha = 0.05

Point estimate = sample proportion = \hat p = x / n = 0.215

This a right (One) tailed test.

The null and alternative hypothesis is,

Ho: p = 0.20

Ha: p > 0.20

Test statistics

z = (\hat p - p_{0} ) / \sqrt{} p_{0}*(1-p_{0}) / n

= ( 0.215 - 0.20) / \sqrt{} (0.20*0.80) / 270

= 0.609

Critical value of  the significance level is α = 0.05, and the critical value for a right-tailed test is

Z_{c} = 1.645

Since it is observed that z = 0.609 < Z_{c} = 1.645 , it is then concluded that the null hypothesis is fail to reject.

Conclusion:

No.